# Continual Learning of Large Language Models: A Comprehensive Survey

### v1

@@ 2.2 Continual Learning
Suppose $T$ sets of observations $\{S_{t}\sim{\mathcal{T}}_{t}\}_{t=1}^{T}$ come in as a sequence, where $\{{\mathcal{T}}_{t}\}_{t=1}^{T}$ denotes the $T$ task distributions . At the learning stage of $t>1$, the sets of observations $\{S_{i}\}_{i=1}^{t-1}$ are not accessible (strong) or partially accessible (relaxed).
Remark.
In early stages of continual learning, works mostly focused on the strong memory constraint [140, 161, 4, 173]; as the research field progresses, more focus was put on relaxing the memory constraint to a small buffer for replay [239, 38, 29, 260]; some modern continual learning works consider the scenario where this constraint is completely discarded but the constraint on the computational budget is present [31, 228, 283].
2.2.1 Three Types of Continual Learning
There are three outstanding types of continual learning scenarios: task-incremental learning (TIL), domain-incremental learning (DIL), and class-incremental learning (CIL). To establish a groundwork for subsequent discussions (as illustrated in Table 3 and Section 6.2), we adhere to the conceptual framework proposed by [282, 139, 288] and offer formal definitions for these three continual learning scenarios.
@@ 2.2.1 Three Types of Continual Learning
$h^{*}$ $=\arg\min_{h}\sum_{t=1}^{T}\mathbb{E}_{({\bm{x}},y)\sim{\mathcal{T}}_{t}}\left[\mathbbm{1}_{h({\bm{x}},t)\neq y}\right].$ (6)
@@
$h^{*}$ $=\arg\min_{h}\sum_{t=1}^{T}\mathbb{E}_{({\bm{x}},y)\sim{\mathcal{D}}_{t}}\left[\mathbbm{1}_{h({\bm{x}})\neq y}\right].$ (7)
@@
$h^{*}$ $=\arg\min_{h}\sum_{t=1}^{T}\mathbb{E}_{({\bm{x}},y)\sim{\mathcal{T}}_{t}}\left[\mathbbm{1}_{h({\bm{x}})\neq(t,y)}\right].$ (8)
Remark.
@@
2.2.2 Techniques of Continual Learning
As outlined in the three definitions provided earlier, the objective of continual learning is to find a hypothesis that minimizes risk across all tasks/domains. Consider domain-incremental learning as an example [260], at $t$-th learning stage, the ideal training objective ${\mathcal{L}}(h)$ is
${\mathcal{L}}(h)$ $_{\text{past domains}}+\underbrace{{\mathcal{L}}_{{\mathcal{D}}_{t}}(h)\vphantom{\sum_{i=1}^{t-1}{\mathcal{L}}_{{\mathcal{D}}_{i}}(h)}}_{\text{current domain}}.$ (9)
The objectives for past domains are often challenging to measure or optimize due to the memory constraints (Definition 2.4). Therefore, the core of designing continual learning algorithms lies in identifying a suitable proxy learning objective for the first term without violating the memory constraint. Existing continual learning techniques can be roughly categorized into 5 groups: (i) replay-based, (ii) regularization-based, (iii) architecture-based, (iv) optimization-based, and (v) architecture-based [61, 288]. Here, we will provide a concise yet comprehensive introduction to the first three categories of continual learning techniques, as they find extensive application in continually learning large language models.
Replay-Based Methods. Replay-based continual learning methods adopt the relaxed memory constraint by keeping a small buffer $\{M_{i}\}_{i=1}^{t-1}$ for each task ${\mathcal{T}}_{i}$ to retain previously observed data examples. Formally, these methods seek to optimize the following empirical training objective:
^$\widehat{{\mathcal{L}}}_{\text{replay}}(h)$ $_{\begin{subarray}{c}\text{proxy for past domains}\end{subarray}}+\underbrace{\widehat{{\mathcal{L}}}_{S_{t}}(h)\vphantom{\sum_{i=1}^{t-1}\widehat{{\mathcal{L}}}_{M_{i}}(h)}}_{\text{current domain}},$ (10)
where ^$\widehat{{\mathcal{L}}}_{S}$ denotes the empirical loss term evaluated on the set of examples $S$. Often regarded as a simplistic solution to continual learning, replay-based methods may theoretically lead to loose generalization bounds [260]. Despite this, they are valued for their simplicity, stability, and high performance, even with a small episodic memory [38, 240]. For instance, DER++ [29] demonstrates consistent performance enhancement by replaying a small set of past examples along with their logits (known as dark experience replay). ESM-ER [250] introduces error sensitivity modulation (ESM) to mitigate abrupt representational drift caused by high-error new examples. A significant focus in replay-based continual learning research is enhancing sample efficiency for buffer maintenance. For instance,  [239] prioritizes exemplar selection based on herding to accurately model class mean throughout class-incremental learning. [361] propose storing low-fidelity examples to achieve memory-efficient exemplar set maintenance. RM (Rainbow Memory) [15] introduces diversity-aware memory updates based on per-sample uncertainty estimation and data augmentation for class-incremental learning.
@@ 2.2.2 Techniques of Continual Learning
^$\widehat{{\mathcal{L}}}_{\text{reg}}(h_{\bm{\theta}})$ $\triangleq\underbrace{\lambda\cdot\left\|{\bm{\theta}}-{\bm{\theta}}_{t-1}\right\|_{\bm{\Sigma}}}_{\begin{subarray}{c}\text{proxy for pastdomains}\end{subarray}}+\underbrace{\widehat{{\mathcal{L}}}_{S_{t}}(h_{\bm{\theta}})\vphantom{}}_{\text{current domain}},$ (11)
@@
where $\|{\bm{v}}\|_{\bm{\Sigma}}={\bm{v}}^{\top}{\bm{\Sigma}}{\bm{v}}$ is the vector norm evaluated on a positive-semi-definite matrix ${\bm{\Sigma}}$, and $\lambda$ is the regularization coefficient, a hyper-parameter introduced to balance the past knowledge retention and current knowledge learning. The matrix ${\bm{\Sigma}}$ introduced is to measure the different level of importance of each parameters and their correlations in retaining the past knowledge. In practice, to reduce computational overhead, diagonal matrices are often designed to encode only the importance of each parameter. For example, Elastic Weight Consolidation (EWC) adopts a Bayesian perspective, using diagonal values from the Fisher Information Matrix (FIM) as an approximation for the Hessian matrix of parameters. This forms a sequential Maximize A Posteriori (MAP) optimization for continual learning [140]. Memory Aware Synapses (MAS) computes parameter importance in an online and unsupervised manner, defining importance by accumulated absolute gradient during training [4]. It is also worth noting that when ${\bm{\Sigma}}={\bm{I}}$ degenerates to an identity matrix, the regularization term simplifies to a basic $l_{2}$-penalty term, evenly penalizing each parameter, which can be surprisingly effective in some cases of continuall learning LLMs [243].
@@
Architecture-Based Methods. Expanding the network architecture dynamically to assimilate new knowledge is deemed the most efficient form of continual learning [300, 299]. This method primarily tackles adaptation challenges and can achieve zero-forgetting when task IDs are available during inference or can be correctly inferred [91, 308]. However, due to the difficulty of task ID inference, architecture expansion is predominantly utilized in task-incremental learning but is scarcely explored in domain-incremental or class-incremental learning. Progressive Neural Networks (PNN) proposes learning laterally connected neurons as new tasks arise, ensuring non-forgetting and enabling transfer of previously learned neurons for future tasks [247]. In conjunction with pre-trained backbone large models like ViT [71], CoLoR [308] trains various low-rank adaptation (LoRA) modules for different tasks. It estimates and stores prototypes for each task and utilizes the natural clustering ability of the pre-trained model during testing to infer task IDs, selecting the corresponding LoRA component for prediction generation. In the domain of continually learning LLMs, architecture expansion has resurged in popularity following the rise of parameter-efficient fine-tuning (PEFT) applied to large models [259, 5, 109, 66, 146, 159], a topic we will delve into shortly [330, 291, 149, 120, 127, 221, 327, 310].
3 Continual Learning Meets Large Language Models: An Overview
@@
Large language models (LLMs) are extensive in various dimensions, including the size of model parameters, pre-training datasets, computational resources, project teams, and development cycles [233, 27, 217, 2, 50, 7, 279, 280]. The substantial scale of LLMs presents notable challenges for development teams, particularly in keeping them updated amidst rapid environmental changes [6, 127, 68, 120, 119]. To illustrate, in 2023, the average daily influx of new tweets posted by users exceeds 500 million [^1] Source: https://www.omnicoreagency.com/twitter-statistics , and training on even a “small” subset of this large volume of data is not affordable. Efficiently and reliably adapting LLMs becomes more critical when considering their cascading impact on downstream applications. Downstream users often lack expertise in collecting and storing large-scale data, maintaining large-scale hardware systems, and training LLMs themselves. Recyclable Tuning [231] is the pioneering study that explicitly outlines the supplier-consumer structure of the modern LLM production pipeline. On the supplier side, the model is continually pre-trained over a sequence of large-scale unlabeled datasets. After every release of the pre-trained model, the consumer needs to utilize the stronger and more up-to-date upstream model for better downstream performance. To enhance the efficiency of fine-tuning for downstream consumers, they initially make several key observations about continually pre-trained LLMs, focusing on mode connectivity and functional similarity. Additionally, they propose reusing the outdated fine-tuned components after a major update of the upstream pre-trained LLM. Building upon the conceptual framework introduced by Recyclable Tuning [231], we present a comprehensive framework for a modern production pipeline encompassing various studies on continual LLM pre-training, adaptation, and deployment, illustrated in Fig. 1. What sets our framework in this survey apart from existing studies [314] is the incorporation of two directions of continuity: vertical continuity and horizontal continuity.
Figure 1: A high-level overview of the modern pipeline for continually pre-training and fine-tuning LLMs, where two dimensions of continuity are described. Vertical Continuity (or Vertical Continual Learning): LLM training can be vertically divided into three stages: (i) Continual Pre-Training (CPT), (ii) Domain-Adaptive Pre-training (DAP), and (iii) Continual Fine-Tuning (CFT). Along the vertical axis, scale of data, scope of tasks, and computational resources, gradually decreases, while the specificity of the LLM is improved towards the final downstream task’s solution. The main focus of vertical continuity is the retention of the LLM’s general knowledge (prevention of vertical forgetting). Horizontal Continuity (or Horizontal Continual Learning): After the LLMs are deployed, the models are continually updated when a new set of data samples becomes available. The primary goal of horizontal continuity is to prevent horizontal forgetting in a long sequence of tasks.
[…]
[175, 120, 127, 119, 6, 68]
[…]
[89, 131, 127, 273, 55, 91, 231, 46, 232]
[…]
[82, 149]
[…]
[177, 209, 323, 158, 87, 326, 221, 49]
[…]
[152, 63, 93, 369, 88, 92, 327, 243, 187]
[…]
[185]
[…]
[310]
[…]
[52]
[…]
[367]
[…]
[20, 311, 331]
[…]
[269, 166, 245, 42, 355, 115]
[…]
[215]
[…]
[134]
[…]
[255, 292, 205, 358, 113]
[…]
[103, 337, 291, 363]
[…]
[162, 96, 108]
[…]
[58, 339, 156]
[…]
[351, 352]
[…]
[165]
[…]
[100, 364, 41]
[…]
[371, 362]
3.1 Vertical Continuity (Vertical Continual Learning)

### v2

@@ 2.2. Continual Learning
Suppose $T$ sets of observations $\{S_{t}\sim{\mathcal{T}}_{t}\}_{t=1}^{T}$ come in as a sequence, where $\{{\mathcal{T}}_{t}\}_{t=1}^{T}$ denotes the $T$ task distributions . At the learning stage $t>1$, the sets of observations $\{S_{i}\}_{i=1}^{t-1}$ are not accessible (strong) or only partially accessible (relaxed).
Remark 0.
In early stages of CL, works mostly focused on the strong memory constraint (kirkpatrick2017overcoming, ; li2017learning, ; aljundi2018memory, ; lomonaco2020rehearsalfree, ); as the research field progresses, more focus was put on relaxing the memory constraint to a small buffer for replay (rebuffi2017icarl, ; chaudhry2019tiny, ; buzzega2020dark, ; shi2024unified, ); some modern CL works completely discard the memory constraint but put focus on the computational budget (cai2021online, ; prabhu2023online, ; verwimp2024continual, ).
2.2.1. Three Types of Continual Learning
There are three outstanding types of continual learning scenarios: task-incremental learning (TIL), domain-incremental learning (DIL), and class-incremental learning (CIL). To establish a groundwork for subsequent discussions (as illustrated in Table 3 and Section 6.2), we adhere to the conceptual framework proposed by (van2022three, ; kim2022theoretical, ; wang2024comprehensive, ) and offer formal definitions for these three continual learning scenarios.
@@ 2.2.1. Three Types of Continual Learning
(5) $h^{*}$ $=\arg\min_{h}\sum_{t=1}^{T}\mathbb{E}_{({\bm{x}},y)\sim{\mathcal{T}}_{t}}\left[\mathbbm{1}_{h({\bm{x}},t)\neq y}\right].$
@@
(6) $h^{*}$ $=\arg\min_{h}\sum_{t=1}^{T}\mathbb{E}_{({\bm{x}},y)\sim{\mathcal{D}}_{t}}\left[\mathbbm{1}_{h({\bm{x}})\neq y}\right].$
@@
(7) $h^{*}$ $=\arg\min_{h}\sum_{t=1}^{T}\mathbb{E}_{({\bm{x}},y)\sim{\mathcal{T}}_{t}}\left[\mathbbm{1}_{h({\bm{x}})\neq(t,y)}\right].$
Remark 0.
@@
2.2.2. Techniques of Continual Learning
The objective of CL is to find a hypothesis that minimizes risk across all tasks/domains. Consider DIL as an example (shi2024unified, ), at $t$-th learning stage, the ideal training objective ${\mathcal{L}}(h)$ is defined as
(8) ${\mathcal{L}}(h)$ $_{\text{past domains}}+\underbrace{{\mathcal{L}}_{{\mathcal{D}}_{t}}(h)\vphantom{\sum_{i=1}^{t-1}{\mathcal{L}}_{{\mathcal{D}}_{i}}(h)}}_{\text{current domain}}.$
The objectives for past domains are often challenging to measure or optimize due to the memory constraints (Definition 2.3). Therefore, the core of designing CL algorithms lies in identifying a proxy learning objective for the first term without violating the memory constraint. Existing CL techniques can be roughly categorized into 5 groups: (i) replay-based, (ii) regularization-based, (iii) architecture-based, (iv) optimization-based, and (v) representation-based (de2021continual, ; wang2024comprehensive, ). Here, we provide a concise yet comprehensive introduction to the first three categories of continual learning techniques, as they find extensive application in continually learning large language models.
Replay-Based Methods. Replay-based methods adopt the relaxed memory constraint by keeping a small buffer of observed data $\{M_{i}\}_{i=1}^{t-1}$ for each task ${\mathcal{T}}_{i}$. Formally, they seek to optimize the following empirical training objective:
(9) ^$\widehat{{\mathcal{L}}}_{\text{replay}}(h)$ $_{\begin{subarray}{c}\text{proxy for past domains}\end{subarray}}+\underbrace{\widehat{{\mathcal{L}}}_{S_{t}}(h)\vphantom{\sum_{i=1}^{t-1}\widehat{{\mathcal{L}}}_{M_{i}}(h)}}_{\text{current domain}},$
where ^$\widehat{{\mathcal{L}}}_{S}$ denotes the empirical loss term evaluated on the set of examples $S$. Often regarded as a simplistic solution to CL, replay-based methods may theoretically lead to loose generalization bounds (shi2024unified, ). Despite this, they are valued for their simplicity, stability, and high performance, even with a small episodic memory (chaudhry2019tiny, ; riemer2018learning, ). For instance, DER++ (buzzega2020dark, ) demonstrates consistent performance enhancement by replaying a small set of past examples along with their logits (known as dark experience replay). ESM-ER (sarfraz2023error, ) introduces error sensitivity modulation (ESM) to mitigate abrupt representational drift caused by high-error new examples. A significant focus in replay-based CL is enhancing sample efficiency for buffer maintenance. For instance,  (rebuffi2017icarl, ) prioritizes exemplar selection based on herding to accurately model class mean throughout class-incremental learning. (zhao2022memory, ) propose storing low-fidelity examples to achieve memory-efficient exemplar set maintenance. RM (Rainbow Memory) (bang2021rainbow, ) introduces diversity-aware memory updates based on per-sample uncertainty estimation and data augmentation for class-incremental learning.
@@ 2.2.2. Techniques of Continual Learning
(10) ^$\widehat{{\mathcal{L}}}_{\text{reg}}(h_{\bm{\theta}})$ $\triangleq\underbrace{\lambda\cdot\left\|{\bm{\theta}}-{\bm{\theta}}_{t-1}\right\|_{\bm{\Sigma}}}_{\begin{subarray}{c}\text{proxy for pastdomains}\end{subarray}}+\underbrace{\widehat{{\mathcal{L}}}_{S_{t}}(h_{\bm{\theta}})\vphantom{}}_{\text{current domain}},$
@@
where $\|{\bm{v}}\|_{\bm{\Sigma}}={\bm{v}}^{\top}{\bm{\Sigma}}{\bm{v}}$ is the vector norm evaluated on a positive-semi-definite matrix ${\bm{\Sigma}}$, and $\lambda$ is the regularization coefficient, a hyper-parameter introduced to balance the past knowledge retention and current knowledge learning. The matrix ${\bm{\Sigma}}$ introduced is to measure the different level of importance of each parameters and their correlations in retaining the past knowledge. In practice, to reduce computational overhead, diagonal matrices are often designed to encode only the importance of each parameter. For example, Elastic Weight Consolidation (EWC) (kirkpatrick2017overcoming, ) adopts a Bayesian perspective, using diagonal values from the Fisher Information Matrix (FIM) as an approximation for the Hessian matrix of parameters. This forms a sequential Maximize A Posteriori (MAP) optimization for continual learning. Memory Aware Synapses (MAS) (aljundi2018memory, ) computes parameter importance in an online and unsupervised manner, defining importance by accumulated absolute gradient during training. It is also worth noting that when ${\bm{\Sigma}}={\bm{I}}$ degenerates to an identity matrix, the regularization term simplifies to a basic $l_{2}$-penalty term, equally penalizing each parameter, which can be surprisingly effective in some cases of continual LLMs (rongali2021continual, ).
@@
Architecture-Based Methods. Expanding the network architecture dynamically to assimilate new knowledge is deemed the most efficient form of CL (wang2022learning, ; wang2022dualprompt, ). This method primarily tackles adaptation challenges and can achieve zero-forgetting when task IDs are available during inference or can be correctly inferred (gururangan2022demix, ; wistuba2023, ). However, due to the difficulty of task ID inference, architecture expansion is predominantly utilized in TIL but is scarcely explored in DIL or CIL. Progressive Neural Networks (PNN) (rusu2016progressive, ) proposes learning laterally connected neurons as new tasks arise, ensuring non-forgetting and enabling transfer of previously learned neurons for future tasks. In conjunction with pre-trained backbone large models like ViT (dosovitskiy2020image, ), CoLoR (wistuba2023, ) trains various low-rank adaptation (LoRA) (hu2021lora, ) modules for different tasks. It estimates and stores prototypes for each task and utilizes the natural clustering ability of the pre-trained model during testing to infer task IDs, selecting the corresponding LoRA component for prediction generation. In the domain of continual LLMs, architecture expansion has resurged in popularity following the rise of parameter-efficient fine-tuning (PEFT) (shazeer2017outrageously, ; aljundi2017expert, ; hu2021lora, ; dettmers2023qlora, ; lester2021power, ; li2021prefix, ), a topic we will delve into shortly (yang2024moral, ; wang2023orthogonal, ; li2024examining, ; jang2022towards, ; jin2022lifelong, ; paul2024ircoder, ; yan2023af, ; wu2024llama, ).
2.2.3. Evaluation Metrics of Continual Learning
In the realm of conventional continual learning, where task streams take the form of classification, many metrics rely on the concept of Accuracy Matrix (lopez2017gradient, ; shi2024unified, ). Extending this notion to the context of continually learning LLMs, we introduce the Performance Matrix ${\bm{P}}\in\mathbb{R}^{T\times T}$, where $T$ represents the total number of training stages. Each entry of ${\bm{P}}$ corresponds to a performance metric evaluated on the models, such as perplexity on pre-training data (jin2022lifelong, ; chen2023lifelong, ; gupta2023continual, ), zero-shot/few-shot evaluation metrics on downstream data without fine-tuning (colombo2024saullm7b, ; wu2023pmc, ; Azerbayev2023LLEMMA, ; deng2023learning, ; nijkamp2022codegen, ; rozière2024code, ), fine-tuned accuracies on downstream tasks (amba2021dynamic, ; qin2023recyclable, ; chen2023lifelong, ; jang2022towards, ), and probing accuracies from fine-tuning add-on components evaluated on downstream tasks (tao2022can, ; luo2023investigating, ; zheng2023learn, ). In ${\bm{P}}$, $P_{i,j}$ denotes the model’s performance after training on task $i$ and evaluating on task $j$. With this Performance Matrix definition, we introduce the primary evaluation protocols widely adopted.
Overall Performance (OP). The Overall Performance (OP) (ke2021achieve, ; zhang2022continual, ; zhang2023copf, ) is a natural extension of the concept of Average Accuracy (lopez2017gradient, ; shi2024unified, ). The OP measured up until training stage $t$ is the average performance of the model trained right after the stage $t$. Denote it as $\operatorname{OP}_{t}$ and we have:
(11) $\operatorname{OP}_{t}$ $\triangleq\tfrac{1}{t}\sum_{i=1}^{t}P_{t,i}.$
As noted in (shi2024unified, ), the OP corresponds to the primary optimization objective defined in Definition 2.4, 2.5, and 2.6. In much of the continual learning literature, once all $T$ tasks are completed, the final $\operatorname{OP}$ ($\operatorname{OP}_{T}$) is reported, with the subscript T often omitted for brevity. In some works, OP is weighted by the importance of tasks $\widetilde{\operatorname{OP}}\triangleq\tfrac{1}{T}\sum_{i=1}^{T}w_{i}P_{t,i}$, where $w_{i}=N_{i}/\sum_{j=1}^{T}N_{j}$ represents the ratio of data.
(12) $F_{t}$ $\triangleq\tfrac{1}{t-1}\sum_{j=1}^{t-1}\left[\max_{l\in[t-1]}\{P_{l,j}-P_{t,j}\}\right].$
Typically, researchers report the average forgetting $F=F_{T}$ at the end of the entire training process. Forgetting quantifies the impact of learning new tasks on previously acquired knowledge. Ideally, a robust continual learning framework should achieve Backward Transfer (BWT), where learning new tasks enhances performance on prior tasks. This enhancement is typically measured by negating the forgetting, thus indicating an improvement in performance on earlier tasks.
(13) $\operatorname{FWT}_{t}$ $\triangleq\tfrac{1}{t-1}\sum_{i=2}^{t}P_{i-1,i}-b_{i},$
where $b_{i}$ is the baseline performance of the model evaluated on task $i$ before undergoing continual learning. Strictly speaking, the definition of $b_{i}$ is not the same as defined in the previous work (lopez2017gradient, ; shi2024unified, ), where it is used to denote the performance of a random initialization of the model.
3. Continual Learning Meets Large Language Models: An Overview
@@
Large language models (LLMs) are extensive in various dimensions, including the size of model parameters, pre-training datasets, computational resources, project teams, and development cycles (radford2019language, ; brown2020language, ; achiam2022chatgpt, ; achiam2023gpt, ; chowdhery2023palm, ; anil2023palm, ; touvron2023llama, ; touvron2023llama2, ). The substantial scale of LLMs presents notable challenges for development teams, particularly in keeping them updated amidst rapid environmental changes (amba2021dynamic, ; jin2022lifelong, ; dhingra2022time, ; jang2022towards, ; jang2022temporalwiki, ). To illustrate, in 2023, the average daily influx of new tweets exceeds 500 million [^1] Source: https://www.omnicoreagency.com/twitter-statistics , and training on even a subset of this large volume of data is unaffordable. Recyclable Tuning (qin2023recyclable, ) is the first work to explicitly outline the supplier-consumer structure in the modern LLM production pipeline. This structure allows us to dissect the challenges of continual LLMs from the perspectives of various roles involved. On the supplier side, the model is continually pre-trained over a sequence of large-scale unlabeled datasets. After every release of the pre-trained model, the consumer needs to utilize the stronger and more up-to-date upstream model for downstream tasks. Compared to the upstream supplier, downstream users often lack capacity of collecting and storing large-scale data, maintaining large-scale hardware systems, and training LLMs themselves. Therefore Recyclable Tuning mainly focuses on efficiently adapting an updated pre-trained LLM to downstream tasks continuously. In this survey, we further present a comprehensive framework for a modern production pipeline encompassing various studies on continual LLM pre-training, adaptation, and deployment (Fig. 1). What sets our framework in this survey apart from existing studies (wu2024continual, ) is the incorporation of two directions of continuity: Vertical Continuity and Horizontal Continuity.
3.1. Vertical Continuity (Vertical Continual Learning)